OpenStudy (saifoo.khan):

Rational Root Theorm: x^3 + 9x^2 +15x-114=0

OpenStudy (anonymous):

ok rational root theorem says that the "rational roots" have to be of the form \[\frac{p}{q}\] where p divides the constant and q divides the leading coefficient.

OpenStudy (saifoo.khan):

Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem.

OpenStudy (anonymous):

since your leading coefficient is 1, that means they will have to be integers that divide 114

OpenStudy (saifoo.khan):

So, how i write it? is that's all the working?

OpenStudy (anonymous):

\[114=2\times 3\times 19\] so you have a bunch of choices: \[\{\pm1,\pm2,\pm3,\pm19,\pm114,\pm6,\pm38,\pm57\]

OpenStudy (saifoo.khan):

which method is that?

OpenStudy (anonymous):

that is finding the possible rational roots

OpenStudy (anonymous):

fundamental theorem doesn't say anything about finding the roots, only says that there is one

myininaya (myininaya):

what study group

OpenStudy (anonymous):

you polynomial must have a root in the complex plane

OpenStudy (anonymous):

at the top it says openstudy is having study groups. i know nothing about it

OpenStudy (anonymous):

because your polynomial is of degree 3 it must have one REAL zero, but it doesn't have to be rational

myininaya (myininaya):

we should ask saifoo about it he knows everything

OpenStudy (anonymous):

factor theorem says if r is a root of p(x) then \[p(x)=(x-r)q(x)\] in other words if you know a zero, you can factor

OpenStudy (saifoo.khan):

here, im blank!! literally.

OpenStudy (anonymous):

any myininaya will have to tell you about descartes rule of sign, because i forget

OpenStudy (saifoo.khan):

Sorry sat, im lost again. which one is: Fundamental Theorem of Algebra,? the Rational Root Theorem,? Factor therom?

OpenStudy (anonymous):

says something about the possible positive or negative zeros having to do with the change of signs of the coefficients of the polynomial

OpenStudy (anonymous):

1) fundamental theorem of algebra: every polynomial has a zero (in the complex plane)

OpenStudy (saifoo.khan):

im trying to figure it out from 5 hours now. :(

OpenStudy (saifoo.khan):

And where is the working for FTA?

OpenStudy (anonymous):

there is no "working" for it. it is just an "existence " theorem. tells you there is one, not how to find it

OpenStudy (saifoo.khan):

tell me the exact words which i should write?

OpenStudy (anonymous):

rational root theorem is as i described.

myininaya (myininaya):

\[f(x)=x^3+9x^2+15x-114\] + + + - there is only 1 change so this means there is one possible real root \[f(-x)=-x^3+9x^2-15x-114\] - + - - 2 changes so this means there are 2 possible or 0 possible negative real roots so there could be 1+ and 2- real roots or there could be 1+ and 0- real roots which means there is 2 complex roots

OpenStudy (anonymous):

"by the fundamental theorem of algebra, \[x^3 + 9x^2 +15x-114=0\] as a solution in the complex plane."

OpenStudy (anonymous):

since the complex zeros come in conjugate pairs, there must be one real zero, since your polynomial has odd degree (it is 3)

OpenStudy (anonymous):

ok so one change in sign of \[p(x)\] means one possible positive root?

myininaya (myininaya):

that one line is suppose to say possible real root

myininaya (myininaya):

positive dop

myininaya (myininaya):

i keep on leaving out the word positive

OpenStudy (anonymous):

and two changes of sign of \[p(-x)\] means 0 or 2 possible negative roots yes?

OpenStudy (saifoo.khan):

Is that correct.

myininaya (myininaya):

yes satellite

OpenStudy (anonymous):

@saifoo i listed the possible rational roots for you

myininaya (myininaya):

we want 3 solutions so we have two possiblilities 1 positive real +2 negative real+0 complex 1 positive real +0 negative real +2 complex

OpenStudy (anonymous):

it is a long list. of course they are only the "possible rational roots" no one guarantees that the roots are rational

OpenStudy (saifoo.khan):

If you guys are free, can you help me out? feel free to decline. i would not mind. can u guys write me on a doc file and send it here? we can divide the word among both.

OpenStudy (saifoo.khan):

I'm really in a terrible situation now!!

OpenStudy (anonymous):

in fact you can check that the one real root is in fact positive and not rational

OpenStudy (saifoo.khan):

work*

OpenStudy (saifoo.khan):

what do you guys say?

myininaya (myininaya):

i don't understand you want us to write what we wrote here but put it in a doc file?

myininaya (myininaya):

why don't you just copy it down

OpenStudy (saifoo.khan):

Yes, this is b/c i can't really copy it from here, it changes it's format.

OpenStudy (saifoo.khan):

Let me do a thing, i will type all this what you wrote in doc file and shot it you you guys.. then just let me know if that's correct or not.

OpenStudy (anonymous):

i cnt help with the prob but i am here so everyones day just got better!!

OpenStudy (anonymous):

give me a minute